OpenStudy (anonymous):

2) Find the ratio of a to b 10/21 = 2a/a+b A) 41/10 B) 10/13 C) 1/13 D) 5/16

OpenStudy (anonymous):

\[\frac{2a}{a+b} = \frac{10}{21}\]

OpenStudy (anonymous):

First take reciprocals both the sides.. \[\frac{a+b}{2a} = \frac{21}{10}\]

OpenStudy (anonymous):

Now separate the terms on left side, with having equal denominators like this: \[\frac{a}{2a} + \frac{b}{2a} = \frac{21}{10}\] You can cancel \(a\) with \(a\).. :)

OpenStudy (anonymous):

\[\frac{1}{2} + \frac{b}{2a} = \frac{21}{10} \implies \frac{b}{2a} = \frac{21}{10} - \frac{1}{2} \implies \frac{b}{2a} = \frac{21-5}{10}\]

OpenStudy (anonymous):

\[\frac{b}{2a} = \frac{16}{10} \implies 2 \times \frac{b}{2a} = 2 \times \frac{16}{10} \implies \frac{b}{a} = \frac{16}{5}\]

OpenStudy (anonymous):

Just take the reciprocals again to get the ratio \(a:b\)